OpenStudy (raiders88):

A science teacher has a supply of 5% hydrochloric acid and a supply of 65% hydrochloric acid (HCl). How much of each solution should the teacher mix together to get 42 mL of 45% HCl for an experiment? (1 point) The teacher will need 14 mL of the 5% solution and 28 mL of the 65% solution. The teacher will need 28 mL of the 5% solution and 14 mL of the 65% solution. The teacher will need 14 mL of the 5% solution and 14 mL of the 65% solution. The teacher will need 28 mL of the 5% solution and 28 mL of the 65% solution.

OpenStudy (here_to_help15):

5% A + 65% B = 45% ( A + B ) A + B = 42 0.05 (42 - B) + 0.65 B = 0.45 x 42 2.10 - 0.05 B + 0.65 B = 18.90 0.60 B = 16.80 B = 28 A = 42 - B = 14 The teacher will need 14 mL of the 5% solution and 28 mL of the 65% solution.

OpenStudy (raiders88):

thx

OpenStudy (here_to_help15):

Np